OpenStudy (anonymous):

Please show work so i can learn:) Solve the systems using substitution 3x + 2y = -4 x + 2y = -8

OpenStudy (anonymous):

So first of all we should know what exactly substitution is. In substitution, we express one variable in terms of the other variable, and then use that in THE OTHER EQUATION. For this question - 3x + 2y = -4 x + 2y = -8 we now write 2y = -3x-4 y = -3x-4/2 Now we will use y = (-3x-4)/2 in x+2y = -8 (DO NOT USE THIS VALUE OF Y IN THE 3X+2Y = -4, U WONT GET ANSWER, I HAD LOST SCORES IN A CERTAIN EXAM BCOZ OF DOING THIS) We have - x + 2(-3x-4)/2 = -8 2/2 = 1, cancelled => x - 3x - 4 = -8 8-4 = 3x-x 4 = 2x, or x = 2 Now you can use. x+2y = -8 2 + 2y = -8 2y = -10, hence y = -5.